Investigating the affect of Regulatory Decisions in a Developmental Model

Pauline C Haddow

CRAB lab

Department of Computer and Information Science The Norwegian University of Science and Technology

Trondheim, Norway, pauline@idi.ntnu.no

Johan Hoye CRAB lab

Department of Computer and Information Science The Norwegian University of Science and Technology

Trondheim, Norway

November 14, 2008